Investigate the linear dependence and independence of the vector X1=(1,2,3), X2=(2,-1,3), X3=(0,1,2), X4=(-3,7,2)

1
Expert's answer
2021-12-12T23:46:46-0500

We are tasked with finding a nonzero solution "\left(a_{1}, a_{2}, a_{3}, a_{4}\right)" to the equation:

 "a_{1} x_{1}+a_{2} x_{2}+a_{3} x_{3}+a_{4} x_{4}=\mathbf{0}"

This is the same as finding a nonzero solution "\left(a_{1}, a_{2}, a_{3}, a_{4}\right)" to the system

 "\left\{\begin{array}{l}

1 a_{1}-3 a_{2}+1 a_{3}+2 a_{2}=0 \\

-1 a_{1}+2 a_{2}+2 a_{3}+3 a_{4}=0 \\

2 a_{1}+1 a_{2}-3 a_{3}+1 a_{4}=0

\end{array}\right."


This is the same as finding a solution "\left[a_{1}, a_{2}, a_{3}, a_{4}\right]^{T}" to the matrix equation

"\left[\begin{array}{cccc}

1 & -3 & 1 & 2 \\

-1 & 2 & 2 & 3 \\

2 & 1 & -3 & 1

\end{array}\right]\left[\begin{array}{l}
a_{1} \\

a_{2} \\

a_{3} \\

a_{4}

\end{array}\right]=\left[\begin{array}{l}

0 \\

0 \\

0

\end{array}\right]"  

In other words, we are trying to find the kernel of the matrix.

Row reducing

 "\operatorname{rref}\left(\left[\begin{array}{cccc}

1 & -3 & 1 & 2 \\

-1 & 2 & 2 & 3 \\

2 & 1 & -3 & 1

\end{array}\right]\right)=\left[\begin{array}{llll}

1 & 0 & 0 & 3 \\

0 & 1 & 0 & 1 \\

0 & 0 & 1 & 2

\end{array}\right]"

 Remembering that row-reducing a system of equations retains the solution-set of the system and reinterpreting this as a system of equations this is the system:

 "\left\{\begin{array}{rrrrr}

1 a_{1} & & & +3 a_{4} & =0 \\

& 1 a_{2} & & +1 a_{4} & =0 \\

& & 1 a_{3} & +2 a_{4} & =0

\end{array}\right."

This tells us that for "\left(a_{1}, a_{2}, a_{3}, a_{4}\right)" to be a solution that "a_{1}=-3 a_{4}, a_{2}=-a_{4}" and "a_{3}=-2 a_{4}"

Picking any number (other than zero) for "a_{4}" , we can then write a linear combination of "x_{1}, \ldots, x_{4}" resulting in zero.